OpenStudy (anonymous):

Quick questions: What is cos multiplied by cos and what is cos divided by cos

Nnesha (nnesha):

let cos = x now what's x times x and what is x divided by x ?

OpenStudy (anonymous):

x^2

OpenStudy (anonymous):

x divided by x equals 1?

Nnesha (nnesha):

right so cos times cos would be \(\large\rm {cos^2 \theta}\)

Nnesha (nnesha):

\[\rm \sin^2 \theta +\cos^2 \theta =1\] this is a special identity which can be written as \[\cos^2 \theta = 1-\sin^2 \theta\] so to verify identity you can replace cos^2 with 1-sin^2

Nnesha (nnesha):

eh but that wasn't your question and yes x/x =1 so cos/cos =1

OpenStudy (anonymous):

and does the same principle apply to sin also?

Nnesha (nnesha):

we are using exponent rules here \[\rm \huge\rm x^m \times x^n =x^{m+n}\] multiply same bases so add their exponents \[\huge\rm \frac{ x^m }{ x^n }=x^{m-n}\] when we divide same bases we should subtract their exponents

Nnesha (nnesha):

and yes right sin is same as sin to the one power \[\frac{ \sin^1 }{ \sin^1 } =\sin^{1-1} = \sin^{0}=\sin \theta\] anything to the zero power equal to one

Nnesha (nnesha):

and when we multiply same bases \[\huge\rm sin^{1} *\sin^{1} =sin^{1+1}=sin^{2} theta \]

OpenStudy (anonymous):

wait, sin/sin=sin?

Nnesha (nnesha):

same thing just like cos

Nnesha (nnesha):

what would you get when you divide same bases ?

Nnesha (nnesha):

like 2/2 ,5/5 y/y